Determine the Power set of 𝐴 × 𝐵 if 𝐴 = { 7, 8 } and 𝐵 = { 𝑎, 𝑏 }


1
Expert's answer
2022-03-02T09:21:10-0500

Solution:

𝐴 = { 7, 8 } and 𝐵 = { 𝑎, 𝑏 }

𝐴 × 𝐵 = {(7,a), (7,b), (8,a), (8,b)}

n(P(𝐴 × 𝐵)) = 42=16

P(𝐴 × 𝐵)={"\phi" , {(7,a)}, {(7,b)}, {(8,a)}, {(8,b)}, {(7,a), (7,b)}, {(7,a),(8,a)}, {(7,a),(8,b)}, {(7,b),(8,a)}, {(7,b),(8,b)}, {(8,a),(8,b)}, {(7,a), (7,b), (8,a)}, {(7,a), (7,b), (8,b)}, {(7,b), (8,a), (8,b)}, {(7,a), (8,a), (8,b)}, {(7,a), (7,b), (8,a), (8,b)} }